OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (parker.goodbar):

the slope of a line is m=(y2-y1)/(x2-x1) where (x1,y1) and (x2,y2) are the two points on the line. So we have x1=0,y1=4 and x2=-8 y2=-1 this gives m=(-1-4)/(-8-0) =5/8

OpenStudy (michele_laino):

here we have to do this computation: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 1 - 4}}{{ - 8 - 0}} = ...?\]